the sides of a number cube are labeled 1 through 6. the probability of randomly rolling a 3 is 1/6. what is the probability of not rolling a 3?

Answer by jim_thompson5910(35256)

P(Not 3) = 1 - P(3)

P(Not 3) = 1 - 1/6

P(Not 3) = 5/6